The life of a compressor manufactured by a company is known to be 200 months on an
average following an exponential distribution .Find the probability that the life of a
compressor is i) less than 200 months ii) between 100 months to 25 years.
1
Expert's answer
2020-04-21T11:45:16-0400

Exponential Distribution Probability 


"F(x;\lambda) = \begin{cases}
 0 &\text{if } x<0 \\
 1-e^{-\lambda x} &\text{if } x\geq0
\end{cases}"

Given "\beta=200, \lambda=\dfrac{1}{\beta}=\dfrac{1}{200}"

 i) less than 200 months


"P(X<200)=F(200;\lambda)=1-e^{-{1 \over 200}( 200)} ="

"=1-e^{-1}\approx0.63212"

ii) between 100 months to 25 years.


"P(100<X<3000)=F(3000;\lambda)-F(100;\lambda)="

"=1-e^{-{1 \over 200}( 3000)} -(1-e^{-{1 \over 200}( 100)} )="

"=e^{-{1 \over 2}}-e^{-15}\approx0.60653066-0.00000031\approx0.60653"